OpenStudy (johan14th):

how to solve for the inverse of f(x)= 3/x-2 y= 3/y-2 ???? what would the next step be?

jimthompson5910 (jim_thompson5910):

is the original function \[\Large f(x) = \frac{3}{x-2}\] OR \[\Large f(x) = \frac{3}{x}-2\]

OpenStudy (johan14th):

Oh sorry about the lack of parenthesis it would be the first one.

jimthompson5910 (jim_thompson5910):

ok thanks

jimthompson5910 (jim_thompson5910):

y= 3/(x-2) x= 3/(y-2) ... swap x and y, now solve for y x(y-2)= 3 x*y-2x = 3 x*y = 3+2x y = (3+2x)/x

jimthompson5910 (jim_thompson5910):

So the inverse is \[\Large f^{-1}(x) = \frac{3+2x}{x}\]

OpenStudy (johan14th):

Oh okay that you that makes a lot of sense now. :)